OpenStudy (anonymous):

I need help with a u-sub \[\int\limits \frac{ dx }{ 1 - x^4 }\]

OpenStudy (baby456):

i got help someone is coming to help you thx for helping me

OpenStudy (anonymous):

Can someone please check me

OpenStudy (anonymous):

\[\frac{ 1 }{ 4 }\ln \left| \frac{ 1+4 }{ 1-x } \right|+\frac{ 1 }{2 }\arctan x + C\]